### v1

# Generalized cryptographic semi-regular sequences: A variant of Fröberg conjecture and a simple complexity estimation for Gröbner basis computation

Abstract
In this paper, we study generalized cryptographic semi-regular sequences, which are expected to generic in the space of homogeneous polynomial sequences on which the coordinate rings have Krull dimension one. We provide an upper-bound on the complexity of the Gröbner basis computation for such a sequence.
1 Introduction
In the theory of polynomial rings over fields, Gröbner bases introduced first by Buchberger [8] provide various methods for investigating important properties of an ideal and its associated algebraic variety. Here, a Gröbner basis is defined as a special kind of its generator set, relative to a monomial ordering. Buchberger also presented the first algorithm (Buchberger’s algorithm) for computing Gröbner bases. Lazard’s formulation in [30] and [31] relates the problem of computing Gröbner bases to the problem of reducing so-called Macaulay matrices. FGLM [19] and Gröbner walk [12] were propsed as algorithms that efficiently convert a given Gröbner basis to a Gröbner basis with respect to another monomial ordering; Hilbert driven [37] can be also applied to such a conversion. Faugère’s proposed his innovative algorithms $F_{4}$ [17] and $F_{5}$ [18] allowing one to quite efficiently compute Gröbner bases, and these are nowadays state-of-the-art algorithms. We refer to [16] for a survey on signature-based algorithms including $F_{5}$. Thanks to Gröbner basis conversion algorithms, it is important to estimate the complexity of computing a Gröbner basis with respect to a specific monomial ordering. In this paper, we consider only the case of graded monomial orderings.
In order to estimate the complexity, it suffices to provide an upper-bound on the so-called solving degree: Roughly speaking, the solving degree is the highest degree of polynomials involved in Gröbner basis computation, see [29, Section 2.2] for rigorous definitions. If the ideal is homogeneous, the solving degree can be equal to the maximal degree of the reduced Gröbner basis with respect to a monomial ordering that one considers, and its upper-bounds have been well-studied. In the case of inhomogeneous ideals, the solving degree can be greater than the maximal degree of the reduced Gröbner basis, whence estimating it directly is difficult in general, see [10] and [29] for details. On the other hand, it is well-known (cf. [27, Prop. 4.3.18]) that a Gröbner basis of an inhomogeneous ideal with respect to a graded monomial ordering is easily obtained from a Gröbner basis of the ideal of homogenized generators. Therefore, the homogeneous case is very important, and we focus on that case. In the following, let $K$ be a field, and let $R=K[x_{1},\ldots,x_{n}]$ (resp. $R^{\prime}=K[x_{0},x_{1},\ldots,x_{n}]$) be the polynomial ring of $n$ (resp. $n+1$) variables over $K$. For a possibly inhomogeneous ideal $I$ of $R$ (resp. $R^{\prime}$), we denote by $\mathrm{max.GB.deg}_{\prec}(I)$ the maximal degree of the elements of the reduced Gröbner basis of $I$ with respect to a monomial ordering $\prec$ on the set of monomials in $R$ (resp. $R^{\prime}$), following [9, Definition 7] (it is also denoted by $\deg(I,\prec)$ in the literature of commutative algebra, see e.g., [24, Definition 4]). If $I$ is generated by $F\subset R$, we may write $\mathrm{max.GB.deg}_{\prec}(F)$ in place of it.
For a homogeneous ideal $I\subset R=K[x_{1},\ldots,x_{n}]$ with a generating set $F=\{f_{1},\ldots,f_{m}\}\subset R$, it is well-known (cf. [2, Proposition 1]) that a Gröbner basis of $I$ with respect to a graded monomial ordering up to degree $d$ can be computed in $O(md\binom{n+d-1}{d}^{\omega})$, where $32\leq\omega<$ 3 is the exponent of matrix multiplication. (We can also obtain another upper-bound $O(m\binom{n+d}{d}^{\omega})$ by the proof of [35, Theorem 1.72].) With this fact, we have in the homogeneous case that the complexity of computing the reduced Gröbner basis of $I$ is upper-bounded by $O(mD\binom{n+D-1}{D}^{\omega})$ (or $O(m\binom{n+D}{D}^{\omega})$) for any integer $D$ with $\mathrm{max.GB.deg}_{\prec}(I)\leq D$. Hence, finding a tighter bound $D$ of $\mathrm{max.GB.deg}_{\prec}(I)$ is important for estimating the complexity of Gröbner basis computation in the homogeneous case. The case on which we focus on is the case where the Krull dimension $\mathrm{Krull.dim}(R/I)$ is equal to or less than one; see e.g., [24, Section 1] for a summary of existing results when $\mathrm{Krull.dim}R/I\geq 2$. In the case where $\mathrm{Krull.dim}(R/I)\leq 1$, Lazard’s bound (called also Macaulay’s bound) is well-known, and it is given by $\sum_{j=1}^{\ell}(d_{j}-1)+1$ with $d_{1}\geq d_{2}\geq\cdots\geq d_{m}$ and $\ell:=\max\{m,n\}$, where we set $d_{j}=\deg(f_{j})$ for $1\leq j\leq m$. When $I$ is in quasi stable position (equivalently $R/\langle I,x_{n}\rangle$ is Artinian, see Lemma 2.1.8 below), we can apply Hashemi-Seiler’s inequality
$\mathrm{max.GB.deg}_{\prec}(I)\leq\mathrm{max}\{\mathrm{hilb}(\langle I,x_{n}\rangle),\mathrm{hilb}(I)\}$ (1.0.1)
for the graded reverse lexicographic ordering $\prec$ with $x_{n}\prec\cdots\prec x_{1}$, where $\mathrm{hilb}(J)$ denotes the Hilbert regularity of a homogeneous ideal $J$ of $R$. Note that $\mathrm{hilb}(\langle I,x_{n}\rangle)$ is equal to the degree of regularity ${d}_{\rm reg}(\langle I,x_{n}\rangle)$ since $R/\langle I,x_{n}\rangle$ is Artinian, and $\mathrm{hilb}(I)$ is equal to the generalized degree of regularity $\widetilde{d}_{\rm reg}(I)$. For the definition of $d_{\rm reg}$ and $d_{\rm reg}$, see Remark 2.1.2 below.
On the other hand, in [29], when $F=\{f_{1},\ldots,f_{m}\}\subset R$ with $m\geq n$ is obtained by homogenizing an affine semi-regular sequence in $K[x_{1},\ldots,x_{n-1}]^{m}$, we improved Lazard’s bound on $\mathrm{max.GB.deg}_{\prec}(\langle F\rangle_{R})$ to $\sum_{j=1}^{n}(d_{j}-1)+1$ with $d_{1}\leq\cdots\leq d_{m}$. Under an additional assumption that $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular (see Definition 2.2.7), we also proved in [29] that the right hand side of (1.0.1) is upper-bounded by the degree of regularity for a cryptographic semi-regular sequence of $m$ polynomials of degrees $d_{1},\ldots,d_{m}$ in $n+1$ variables.
In this paper, we shall prove that, if $K$ is large enough, the inequality (1.0.1) always holds after a linear coordinate change. Then, we also estimate the complexity of computing a Gröbner basis for $I$, assuming that $I$ is generated by a generalized cryptographic semi-regular sequence. Such a sequence is expected to be generic (see Definition 2.3.1 for definition), analogously to Fröberg’s conjecture [20]. Our results are summarized into the following theorem and proposition:
Theorem 1.
Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $I$ be a proper homogeneous ideal of $R$. Let $F=\{f_{1},\ldots,f_{m}\}$ be a subset of $R$ generating $I$, and put $d_{j}:=\deg(f_{j})$ (total degree) for $1\leq j\leq m$. Assume that $I$ has finitely many projective zeros over $\overline{K}$, and suppose that the cardinality of $K$ is greater than the number of projective zeros over $\overline{K}$ of $I$. Choose an arbitrary linear form $\ell$ in $R$ such that it does not vanish at any projective zero over $\overline{K}$ of $I$ (such an $\ell$ always exists by our assumption $\mathrm{Krull.dim}(R/I)\leq 1$). Then, the following (1) holds:
  1. (1)
    There exists a linear transformation $\sigma$ of variables $x_{1},\ldots,x_{n}$ over $K$ sending $\ell$ to $x_{n}$ such that
    ${\rm max.GB.deg}_{\prec}(I^{\sigma})\leq\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}$ (1.0.2)
    for the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$, where $I^{\sigma}$ is the image of $I$ by $\sigma$.
Moreover, if $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular, then following (2) and (3) also hold:
  1. (2)
    We have
    $\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}\leq D^{(n,m)},$ (1.0.3)
    where
    $D^{(n,m)}:=\left\{\!\begin{array}[]{cc}\displaystyle{\deg\left(\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right]\right)}+1&(m\geq n),\\ \displaystyle\sum_{i=1}^{n-1}(d_{i}-1)+1&(m=n-1).\end{array}\right.$ (1.0.4)
  2. (3)
    The complexity of computing a Gröbner basis of $I^{\sigma}$ with respect to $\prec$ is upper-bounded by
    $O\left(m\binom{n+D^{(n,m)}-1}{D^{(n,m)}}^{\omega}\right),$ (1.0.5)
    where $32\leq\omega<$ 3 is the exponent of matrix multiplication.
Proposition 1.
Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $\prec$ be the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$. Let $I$ be a proper homogeneous ideal of $R$. Assume that the Krull dimension of $R/I$ is $\leq 1$ and $R/\langle I,x_{n}\rangle$ is Artinian. If $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic, then the inequality (1.0.1) becomes an equality. Hence, the inequality (1.0.2) in Theorem 1 (2) also becomes an equality.
Theorem 1 can be viewed as a generalization of [29, Proposition 1 and Corollary 1]. The bound $D_{\rm new}$ given in (1.0.4) coincides with Lazard’s bound when $m\leq n$, while it can be sharper than Lazard’s one when $m>n$. We also note that, the complexity bound (1.0.5) is better than the usual bound $O(mD\binom{n+D-1}{D}^{\omega})$ (see e.g., [2, Proposition 1], [36, Equation (24)]) and than $O(m\binom{n+D}{D}^{\omega})$ (which comes from the proof of [35, Theorem 1.72]) for an arbitrary $D$ with $D\geq\mathrm{max.GB.deg}_{\prec}(F)$.
Our assumption that a homogeneous polynomial sequence is generalized cryptographic semi-regular could be useful to analyze the security of multivariate cryptosystems and so on (cf. [29, Section 4.3]).
Remark 1 (Inhomogeneous case).
By Theorem 1, we can also estimate the complexity of computing a Gröbner basis for an inhomogeneous polynomial sequence defining a zero-dimensional ideal, as in [29, Section 4.3]. Indeed, for a subset $F=\{f_{1},\ldots,f_{m}\}$ of $R=K[x_{1},\ldots,x_{n}]$ containing at least one inhomogeneous polynomial, a Gröbner basis of $\langle F\rangle_{R}$ is obtained by reducing the associated degree-$D$ Macaulay matrix at $D=\mathrm{max.GB.deg}_{\prec^{h}}(I)$ with $I=\langle f_{1}^{h},\ldots,f_{m}^{h}\rangle\subset R^{\prime}:=R[y]$, see [9, Theorem 7]. Here, each $f_{j}^{h}$ denotes the homogenization of $f_{j}$ by an extra variable $y$, and $\prec^{h}$ is the homogenization of $\prec$, see e.g., [27, Section 3.1.1] for definition. The complexity of reducing the Macaulay matrix can be easily estimated as $O(m\binom{n+D}{D}^{\omega})$, without assuming $D=O(n)$. Suppose the following three conditions:
  • The Krull dimension of $R^{\prime}/\langle f_{1}^{h},\ldots,f_{m}^{h},y\rangle$ is zero, namely the Krull dimension of $R/\langle f_{1}^{\rm top},\ldots,f_{m}^{\rm top}\rangle$ is zero, where $f_{j}^{\rm top}:=(f_{j}^{h})|_{y=0}$ for $1\leq j\leq m$. (Equivalently these two graded rings are both Artinian.)
  • The sequence $(f_{1}^{h},\ldots,f_{m}^{h},y)\in(R^{\prime})^{m+1}$ is cryptographic semi-regular, equivalently $(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})\in R^{m}$ is cryptographic semi-regular.
  • The sequence $(f_{1}^{h},\ldots,f_{m}^{h})\in(R^{\prime})^{m+1}$ is generalized cryptographic semi-regular.
Then, by the first assumption, it follows from Theorem 1 (1) that the inequality (1.0.2) holds for $I=\langle F^{h}\rangle_{R^{\prime}}$ and $\ell=y$, where we replace $R$ by $R^{\prime}$ and set $\sigma$ to be the identity. Moreover, under the second and third assumptions, we also have the complexity bound
$O\left(m\binom{n+D^{(n+1,m)}}{D^{(n+1,m)}}^{\omega}\right)$ (1.0.6)
by Theorem 1 (2). Note that this upper-bound is heuristically found in [21].
2 Preliminaries
Let $K$ be a field, and let $\overline{K}$ denote its algebraic closure. Let $n$ be a natural number, and let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over $K$. For $d\geq 0$, we denote by $R_{d}$ the homogeneous component of $R$ of degree $d$. The Krull dimension of a commutative ring $A$ is denoted by $\mathrm{Krull.dim}(A)$. For ideals $I$ and $J$ of $R$, their ideal quotient is defined as $(I:J)=\{f\in R\mid fJ\subset I\}$. The saturation of $I$ with respect to $J$ is $(I:J^{\infty}):=\{f\in R\mid fJ^{s}\subset I\mbox{ for some $s\geq 1$}\}$. In particular, the saturation of a homogeneous ideal $I$ with respect to the maximal homogeneous ideal $\mathfrak{m}=\langle x_{1},\ldots,x_{n}\rangle_{R}$ is denoted by $I^{\rm sat}=(I:\mathfrak{m}^{\infty})$, which is also homogeneous.
2.1 Hilbert regularity
Let $I$ be a homogeneous ideal of $R$. Let ${\rm HP}_{R/I}\in\mathbb{Q}[z]$ denote the Hilbert polynomial of $R/I$, namely there exists a non-negative integer $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for any $d$ with $d\geq d_{0}$, where $\mathrm{HF}$ is the Hilbert function.
Definition 2.1.1 (Hilbert regularity (or index of regularity [13, Chapter IX, Section 3])).
The smallest $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HP}_{R/I}(d)$ for any $d$ with $d\geq d_{0}$ is called the Hilbert regularity of $I$, denoted by ${\rm hilb}(I)$, or is called the index of regularity of $I$, denoted by $i_{\rm reg}(I)$.
Throughout the rest of this subsection, we set $r:=\mathrm{Krull.dim}(R/I)$.
Remark 2.1.2.
It is well-known that the following conditions are all equivalent:
  1. (1)
    The Krull dimension $r$ of $R/I$ is equal to $1$ (resp. 00).
  2. (2)
    The projective algebraic set $V_{\overline{K}}(I)$ is a finite set (resp. an empty set).
  3. (3)
    The Hilbert polynomial ${\rm HP}_{R/I}$ is a positive constant (resp. the zero).
  4. (4)
    There exists a non-negative integer $d_{0}$ such that ${\rm HF}_{R/I}(d)={\rm HF}_{R/I}(d_{0})$ (resp. ${\rm HF}_{R/I}(d)=0$) for any $d$ with $d\geq d_{0}$.
In this case, it is straightforward that
$\mathrm{hilb}(I)$ $=$ $\min\{d_{0}:{\rm HF}_{R/I}(d)={\rm HF}_{R/I}(d_{0})\mbox{ for any$d$ with $d\geq d_{0}$}\}$
$=$ $\min\{d_{0}:\dim_{K}(R/I)_{d}=\dim_{K}(R/I)_{d_{0}}\mbox{ for any$d$ with $d\geq d_{0}$}\}.$
The middle and the right hand sides are called the generalized degree of regularity of $I$ for $r\leq 1$, denoted by $\widetilde{d}_{\rm reg}(I)$ (cf. [29, Sect. 2.3]): When $r=0$, these are equal to the degree of regularity $d_{\rm reg}(I):=\min\{d:R_{d}=I_{d}\}$. Namely, we have
$\mathrm{hilb}(I)=\left\{\begin{array}[]{cll}\widetilde{d}_{\rm reg}(I)&&(r=1),\\ \widetilde{d}_{\rm reg}(I)&=d_{\rm reg}(I)&(r=0).\end{array}\right.$ (2.1.1)
See [7] for bounds on ${d}_{\rm reg}(I)$.
Although we mainly treat the case where $r\leq 1$ in this paper, we distinguish $\mathrm{hil}(I)$ and $\widetilde{d}_{\rm reg}(I)$ for the compatibility with the condition that a sequence of homogeneous polynomials in $R$ is regular (cf. Lemma 2.2.9 below).
Remark 2.1.3.
We also note that ${\rm HS}_{R/I}$ is written as
${\rm HS}_{R/I}(z)=\dfrac{h_{R/I}(z)}{(1-z)^{r}}$
for some polynomial $h_{R/I}(z)\in\mathbb{Z}[z]$, which is called the $h$-polynomial of $R/I$. Then, we have ${\rm hilb}(I)=\deg h_{R/I}-r+1$, see [6, Prop. 4.1.12].
Moreover, it follows from Grothendieck-Serre’s formula that
${\rm hilb}(I)\leq\mathrm{reg}(I)=\mathrm{reg}(R/I)+1,$
where ‘${\rm reg}$’ means Castelnuovo-Mumford regularity, see [15, §20.5] for the definition. The equality holds if $R/I$ is Artinian, i.e., the Krull dimension of $R/I$ is zero.
Definition 2.1.4 (cf. [24, Section 3]).
We say that $I$ is in Noether position with respect to the variables $x_{1},\ldots,x_{n-r}$, where $r=\mathrm{Krull.dim}(R/I)$, if the injective ring extension $K[x_{n-r+1},\ldots,x_{n}]\to R/I$ is integral.
Lemma 2.1.5 ([5, Lemma 4.1]; see also [32, Proposition 5.4]).
Let $I$ be a homogeneous ideal of $R$, and let $\prec$ be the reverse lexicographic order on $R$ with $x_{n}\prec\cdots\prec x_{1}$. Then, the following are equivalent:
  1. (1)
    $I$ is in Noether position with respect to the variables $x_{1},\ldots,x_{n-r}$.
  2. (2)
    For each $i$ with $1\leq i\leq n-r$, there exists $r_{i}\geq 0$ such that $x_{i}^{r_{i}}\in\langle\mathrm{LM}_{\prec}(I)\rangle$.
  3. (3)
    $\mathrm{Krull.dim}(R/\langle I,x_{n-r+1},\ldots,x_{n}\rangle)=0$.
  4. (4)
    $\mathrm{Krull.dim}(R/\langle\langle\mathrm{LM}_{\prec}(I)\rangle,x_{n-r+1},\ldots,x_{n}\rangle)=0$.
In the proof of [5, Lemma 4.1], the authors first proved the equivalence of the conditions (2), (3), and (4), and the implication (1) $\Rightarrow$ (2). The non-trivial part is that (2), (3), and (4) imply (1), which follows from [32, Proposition 5.4].
Here, we also recall the definition introduced by Hashemi [23] that a homogeneous ideal is in strong Noether position.
Definition 2.1.6 ([23, Theorem 2.19]).
A homogeneous ideal $I$ of $R$ is said to be in strong Noether position (SNP) with respect to the ordered variables $x_{1},\ldots,x_{n}$ if we have $I^{\rm sat}=(I:x_{n}^{\infty})$ and
$\langle I,x_{n-i+1},\ldots,x_{n}\rangle^{\rm sat}=(\langle I,x_{n-i+1},\ldots,x_{n}\rangle:x_{n-i}^{\infty})$
for any $i$ with $11\leq i\leq r-$ 1.
For equivalent definitions, see [23, Section 2] for details. Note that $I$ is in SNP if and only if $\langle\mathrm{LM}_{\prec}(I)\rangle$ is in SNP. Hashemi also proved the following inequality:
Theorem 2.1.7 ([23, Proposition 4.7 and Theorem 4.17]).
If $I$ is in SNP, then $\mathrm{reg}(I)$ is equal to
$\overline{\rm hilb}(I):=\max\{\mathrm{hilb}(I)\}\cup\{\mathrm{hilb}(\langle I,x_{n-i+1},\ldots,x_{n}\rangle)\mid 1\leq i\leq{\rm Krull.dim}(R/I)\},$
which is called the stabilized Hilbert regularity of $I$, and moreover
$\mathrm{max.GB.deg}_{\prec}(I)\leq\overline{\rm hilb}(I)$
for the graded lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
Lemma 2.1.8 ([24, Lemma 23] and [25, Proposition 3.2 (v)]).
If $r=1$, then the following are all equivalent:
  1. (1)
    $I$ is in Noether position with respect to $x_{1},\ldots,x_{n-1}$ (or one of the equivalent conditions given in Lemma 2.1.5 holds).
  2. (2)
    $I$ is in quasi stable position with respect to the ordered variables $x_{1},\ldots,x_{n}$ (see e.g., [24, Definition 6] for definition).
  3. (3)
    $I$ is in strong Noether position.
Proof.
The equivalence of (1) and (2) is proved in [24, Lemma 23], and that of (2) and (3) follows from [25, Proposition 3.2 (v)]. ∎
By Theorem 2.1.7 and Lemma 2.1.8, we obtain the following theorem:
Theorem 2.1.9 ([24, Proof of Theorem 30]).
Assume that $r=1$. If $R/\langle I,x_{n}\rangle$ is Artinian (or one of the equivalent conditions given in Lemma 2.1.8 holds), then we have the inequality (1.0.1) for the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$.
2.2 Generalized semi-regular sequences
In this subsection, we recall the definition of cryptographic semi-regular sequences and that of generalized ones. Throughout this subsection, let $f_{1},\ldots,f_{m}\in R$ be homogeneous polynomials of positive degrees $d_{1},\ldots,d_{m}$, and put $I=\langle f_{1},\ldots,f_{m}\rangle$.
Theorem 2.2.1 (cf. [14, Theorem 1]).
For each $d$ with $d\geq\mathrm{max}\{d_{i}:1\leq i\leq m\}$, the following are equivalent:
  1. (1)
    For each $i$ with $1\leq i\leq m$, if a homogeneous polynomial $g\in R$ satisfies $gf_{i}\in\langle f_{1},\ldots,f_{i-1}\rangle_{R}$ and $\mathrm{deg}(gf_{i})<d$, then we have $g\in\langle f_{1},\ldots,f_{i-1}\rangle_{R}$, namely, the multiplication-by-$f_{i}$ map $(R/\langle f_{1},\ldots,f_{i-1}\rangle)_{t-d_{i}}\longrightarrow(R/\langle f_{1},\ldots,f_{i-1}\rangle)_{t}$ is injective for any $t$ with $d_{i}\leq t<d$.
  2. (2)
    We have
    ${\rm HS}_{R/I}(z)\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{d}},$ (2.2.1)
    where $\mathrm{HS}$ denotes the Hilbert series.
  3. (3)
    We have $H_{1}(K_{\bullet}(f_{1},\ldots,f_{m}))_{\leq d-1}=0$, where $K_{\bullet}(f_{1},\ldots,f_{m})$ denotes the Koszul complex on the sequence $(f_{1},\ldots,f_{m})$.
Definition 2.2.2 ([1, Definition 3]; see also [14, Definition 1]).
For each $d$ with $d\geq\mathrm{max}\{d_{i}:1\leq i\leq m\}$, a sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is said to be $d$-regular if the equivalent conditions in Theorem 2.2.1 hold.
Here, we define a cryptographic semi-regular sequence:
Definition 2.2.3 ([1, Definition 5], [3, Definition 5]; see also [14, Section 2]).
A sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is said to be cryptographic semi-regular if it is $d_{\rm reg}(I)$-regular.
The Hilbert series of $R/I$ for a cryptographic semi-regular sequence $(f_{1},\ldots,f_{m})$ is easily computed, by the following Proposition 2.2.4:
Proposition 2.2.4 ([14, Proposition 1 (d)]; see also [3, Proposition 6]).
With the same notation as in Theorem 2.2.1, a sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is cryptograpic semi-regular if and only if
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)=\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right],$ (2.2.2)
where []delimited-[][$\cdot$][ ⋅ ] means truncating a formal power series over $\mathbb{Z}$ after the last consecutive positive coefficient. (Note that, in this case, we have $d_{\rm reg}(I)=\deg({\rm HS}_{R/I}(z))+1$.)
The above ‘standard’ form (2.2.2) of Hilbert series was introduced by Fröberg in his celebrated paper [20]. We will recall his famous conjecture (Conjecture 2.3.3 below) in the next subsection.
Here, we recall the following lemma:
Lemma 2.2.5 ([20, b) on page 121]).
With notation as above, we have
${\rm HS}_{R/I}(z)\geq^{\rm lex}\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right]$
in the lexicographic sense, i.e., the first coefficient that differs, if any, is greater on the left hand side, see [20, page 118].
Remark 2.2.6.
For any non-negative integer $d$, it follows from Lemma 2.2.5 that
$\left({\rm HS}_{R/I}(z)\bmod{z^{d}}\right)\geq^{\rm lex}\left(\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\bmod{z^{d}}\right).$
To deal with the case where $R/I$ has Krull dimension one, we extend the notion of cryptgraphic semi-regular as follows:
Definition 2.2.7 ([29, Definition 2.3.3]).
A sequence $(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$ is said to be generalized cryptographic semi-regular if it is $\widetilde{d}_{\rm reg}(I)$-regular.
Lemma 2.2.8.
If the sequence $\bm{F}=(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular, then we have
$\widetilde{d}_{\rm reg}(I)\leq{\deg\left(\left[\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\right]\right)}+1.$ (2.2.3)
Proof.
By our assumption that $\bm{F}$ is generalized cryptographic semi-regular, the equality (2.2.1) holds for $d=\widetilde{d}_{\rm reg}(I)$. For any $d^{\prime}$ with $d^{\prime}<\widetilde{d}_{\rm reg}(I)$, we have ${\rm HF}_{R/I}(d^{\prime})>0$, whence the $z^{d^{\prime}}$-coefficient of $\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}$ is also positive. Thus, the inequality (2.2.3) holds. ∎
For a sequence $\bm{F}=(f_{1},\ldots,f_{m})$ of homogeneous polynomials in $R\smallsetminus K$, we consider the following four conditions:
$\bm{F}$ is regular, equivalently
${\rm HS}_{R/I}(z)=\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}.$
$\bm{F}$ is cryptographic semi-regular (i.e., $d_{\rm reg}(I)$-regular).
$\bm{F}$ is generalized cryptographic semi-regular (i.e., $\widetilde{d}_{\rm reg}(I)$-regular).
$\bm{F}$ is $\mathrm{hilb}(I)$-regular.
Lemma 2.2.9.
For conditions (A), (B), (C), and (D) as defined above, there is an implication:
$(A)\Rightarrow(B)\Rightarrow(C)\Rightarrow(D),\mbox{ but }(D)\overset{r\geq 2}{\not\Rightarrow}(C)\overset{r=1}{\not\Rightarrow}(B)\overset{m>n}{\not\Rightarrow}(A).$
Namely, (A) and (B) are equivalent for $m\leq n$, (B) and (C) are equivalent for $r\neq 1$, (C) and (D) are equivalent for $r\leq 1$, and (A), (B), (C), and (D) are all equivalent for $m=n$ and $r=0$.
Proof.
It is clear that $(A)\Rightarrow(B)\Rightarrow(C)\Rightarrow(D)$. When $r\leq 1$, it follows from (2.1.1) that $(D)\Rightarrow(C)$ holds. Also when $r\neq 1$, we have $d_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I)$, whence $(C)\Rightarrow(B)$ holds. For $m\leq n$, Proposition 2.2.4 implies $(B)\Rightarrow(C)$. ∎
2.3 Genericness and a variant of Fröberg’s conjecture
Given $m$ non-negative integers $d_{1},\ldots,d_{m}$, let $V(n,m,d_{1},\ldots,d_{m})\!=\!R_{d_{1}}\times\cdots\times R_{d_{m}}$, which is topologically identified with the affine space $\mathbb{A}^{N_{1}}(K)\times\dots\times\mathbb{A}^{N_{m}}(K)$ of dimension $\sum_{j=1}^{m}N_{j}$ with $N_{j}:=\binom{n+d_{j}-1}{d_{j}}$. Here the identification from $R_{d}$ to $\mathbb{A}^{N}$, where $N=\binom{n+d-1}{d}$, is given as follows: Letting $\{t_{1},\ldots,t_{N}\}$ be the set of all monomials in $R$ of degree $d$, each polynomial $f=\sum_{i=1}^{N}c_{i}t_{i}$ in $R_{d}$ is identified with its coefficients vector $(c_{1},\ldots,c_{N})$. We write simply $V$ for $V(n,m,d_{1},\ldots,d_{m})$ when we fix the parameter $(n,m,d_{1},\ldots,d_{m})$.
Definition 2.3.1 (cf. [34, Section 1]).
For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, a property $\mathcal{P}$ for $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is said to be generic (or, we say that a generic sequence in $V$ satisfies $\mathcal{P}$) if it holds on a non-empty Zariski open set in $V$, namely there exists a non-empty Zariski open set $U$ in $V$ such that
$U\subset\{\bm{F}=(f_{1},\ldots,f_{m})\in V:\mbox{$\bm{F}$ satisfies $\mathcal{P}$}\}.$
Remark 2.3.2.
Another definition of genericness is the following (cf. [20, page 119]): For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, a property $\mathcal{P}$ for $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is said to be generic over $K$ if it holds for any $\bm{F}=(f_{1},\ldots,f_{m})\in V$ such that the coefficients of $f_{1},\ldots,f_{m}$ are algebraically independent over the prime field $L$ of $K$, that is, $L=\mathbb{Q}$ or $\mathbb{F}_{p}$ for some prime number $p$. It can be checked that the two definitions of genericness are equivalent.
Here, we recall Fröberg’s conjecture:
Conjecture 2.3.3 (Fröberg’s conjecture [20]).
Assume that $K$ is an infinite field. For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}=(f_{1},\ldots,f_{m})\in V$ satisfies (2.2.2) (i.e., $\bm{F}$ is cryptographic semi-regular) is generic. In other words, a generic sequence $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is cryptographic semi-regular.
See Pardue’s paper [34] for some conjectures equivalent to Fröberg’s one (Conjecture 2.3.3). He also proved that Moreno-Socías conjecture [33] is stronger than Conjecture 2.3.3, see [34, Theorem 2] for a proof.
Here, as a variant of Conjecture 2.3.3, we raise the following conjecture (cf. [29, Conjecture 4.3.4] as its strict version):
Conjecture 2.3.4.
Assume that $K$ is an infinite field. For given $n$, $m$, and $(d_{1},\ldots,d_{m})$, the property that $\bm{F}=(f_{1},\ldots,f_{m})\in V$ is generalized cryptographic semi-regular is generic.
We shall prove Conjectures 2.3.3 and 2.3.4 are equivalent to each other. For this, we first prove the following lemma:
Lemma 2.3.5.
For $m\geq n$, the property that $R/\langle f_{1},\ldots,f_{m}\rangle$ is Artinian is generic.
Proof.
The case $m=n$ is well-known to hold. Assume $m>n$, fix $n$, $m$, and $(d_{1},\ldots,d_{m})$, and put
$\Lambda^{\rm Art}:=\{\bm{F}=(f_{1},\ldots,f_{m})\in V:\mbox{$R/\langle f_{1},\ldots,f_{m}\rangle$ is Artinian}\}.$
Consider a natural projection
$V\to V_{n}:=\mathbb{A}^{N_{1}}(K)\times\dots\times\mathbb{A}^{N_{n}}(K)\ ;\ (f_{1},\ldots,f_{n},f_{n+1},\ldots,f_{m})\mapsto(f_{1},\ldots,f_{n}).$
Then we have $\pi^{-1}(\Lambda_{n}^{\rm Art})\subset\Lambda^{\rm Art}$, where
$\Lambda_{n}^{\rm Art}:=\{\bm{F}=(f_{1},\ldots,f_{n})\in V:\mbox{$R/\langle f_{1},\ldots,f_{n}\rangle$ is Artinian}\}$
is Zariski-open in $V_{n}$. Since $\pi$ is continuous, $\pi^{-1}(\Lambda_{n}^{\rm Art})$ is Zariski-open in $V$, whence the assertion holds. ∎
Proposition 2.3.6.
Assume that $K$ is an infinite field. For $m\geq n$, our conjecture (Conjecture 2.3.4) is equivalent to Fröberg’s conjecture (Conjecture 2.3.3).
Proof.
Fixed $n$, $m$, and $(d_{1},\ldots,d_{m})$, we set
[…]
so that $\Lambda^{\rm csr}=\Lambda^{\rm gcsr}\cap\Lambda^{\rm Art}$.
Assume that $(f_{1},\ldots,f_{m})\in V$ is generically generalized cryptograhic semi-regular, i.e., there exists a non-empty Zariski-open set $U_{1}\subset V$ such that $U_{1}\subset\Lambda^{\rm gcsr}$. Recall from Lemma 2.3.5 that $(f_{1},\ldots,f_{m})\in V$ generically has the coordinate ring $R/\langle f_{1},\ldots,f_{m}\rangle$ that is Artinian, whence there exists a non-empty Zariski-open set $U_{2}\subset V$ such that $U_{2}\subset\Lambda^{\rm Art}$. Since both $U_{1}$ and $U_{2}$ are dense, we have $\emptyset\neq U_{1}\cap U_{2}\subset\Lambda^{\rm gcsr}\cap\Lambda^{\rm Art}=\Lambda^{\rm csr}$. Hence, $(f_{1},\ldots,f_{m})\in V$ is cryptographic semi-regular on the non-empty Zariski open set $U_{1}\cap U_{2}$ of $V$, as desired.
The converse is obvious by $\Lambda^{\rm csr}\subset\Lambda^{\rm gcsr}$. ∎
Despite of the equivalence between two conjectures, those cannot guarantee the genericness of $\Lambda^{\rm gscr}\cap(V\smallsetminus\Lambda^{\rm Art})$ with respect to the induced topology of $V\smallsetminus\Lambda^{\rm Art}$. Once the genericness holds for the case, we have the following.
Remark 2.3.7.
If the subset $\Lambda^{\rm gcsr}\cap(V\smallsetminus\Lambda^{\rm Art})$ contains a non-empty Zariski-open set in the subspace $V\smallsetminus\Lambda^{\rm Art}$ equipped with the induced topology, then ‘almost’ all sequences $(f_{1},\ldots,f_{m})$ such that $R/\langle f_{1},\ldots,f_{m}\rangle$ has Krull dimension $\geq 1$ are expected to be generalized cryptographic semi-regular.
As the first discussion about the genericness of $\Lambda^{\rm gcsr}\cap(V\smallsetminus\Lambda^{\rm Art})$, we consider the following sub case: Let $Z$ be the non-empty Zariski closed set in $V=V(n,m,d_{1},\ldots,d_{m})$ defined by the condition that the $x_{n}^{d_{i}}$-coefficient of $f_{i}$ is zero for any $i$ with $1\leq i\leq m$. Then, for any element $\bm{F}=(f_{1},\ldots,f_{m})\in Z$, the homogeneous ideal $\langle f_{1},\ldots,f_{m}\rangle_{R}$ has a projective root $(0:\cdots:0:1)$, whence $R/\langle f_{1},\ldots,f_{m}\rangle$ is not Artinian. Namely, we have $Z\subset V\smallsetminus\Lambda^{\rm Art}$.
Lemma 2.3.8.
Let $\bm{G}=(g_{1},\ldots,g_{m})\in Z$, and assume that the other coefficients of $g_{1},\ldots,g_{m}$ are algebraically independent over the prime field of $K$. Then, for any $\bm{F}=(f_{1},\ldots,f_{m})\in Z$, we have the coefficient-wise inequality
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}(z)\geq{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}(z).$
Therefore, if there exists a generalized cryptographic semi-regular sequence $\bm{F}=(f_{1},\ldots,f_{m})$ in $Z$ with $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$, then $\bm{G}=(g_{1},\ldots,g_{m})$ is generalized cryptographic semi-regular, i.e.,
${\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}(z)\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D^{\prime}}}$
for $D^{\prime}=\widetilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})$, and $\dim_{K}(R/\langle g_{1},\ldots,g_{m}\rangle)_{D^{\prime}}=1$.
Proof.
The first half can be proved in a way similar to the proof of [20, Lemma 1]. To prove the second one, we also use Lemma 2.2.5 (cf. [20]) for a sequence $\bm{H}=(h_{1},\ldots,h_{m})$ of generic forms with the same parameters $n,m,d_{1},\ldots,d_{m}$. Then, we have ${\rm HS}_{R/\langle\bm{G}\rangle}\geq{\rm HS}_{R/\langle\bm{H}\rangle}$ by [20, Lemma 1] and thus,
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\geq{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\geq{\rm HS}_{R/\langle h_{1},\ldots,h_{m}\rangle}\geq^{\rm lex}\left[\frac{\prod_{i=1}^{m}(1-z^{d_{i}})}{(1-z)^{n}}\right],$ (2.3.1)
where $\geq^{\rm lex}$ denotes the lexicographical inequality. As $\bm{F}$ is a generalized cryptographic semi-regular sequence, we also have
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D}}$ (2.3.2)
for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$. Combining (2.3.1) and (2.3.2), we have
${\rm HS}_{R/\langle f_{1},\ldots,f_{m}\rangle}\equiv{\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D}}.$
Meanwhile, it follows from the condition $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ that we obtain $\dim_{K}(R/\langle g_{1},\ldots,g_{m}\rangle)_{D}=1$ and $\tilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})\leq D$. Hence we have
${\rm HS}_{R/\langle g_{1},\ldots,g_{m}\rangle}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\pmod{z^{D^{\prime}}}$
for $D^{\prime}=\tilde{d}_{\rm reg}(\langle g_{1},\ldots,g_{m}\rangle_{R})$. ∎
From Lemma 2.3.8 and Remark 2.3.2, if there exists a generalized cryptographic semi-regular sequence $\bm{F}=(f_{1},\ldots,f_{m})$ with $\dim_{K}(R/\langle f_{1},\ldots,f_{m}\rangle)_{D}=1$ for $D=\tilde{d}_{\rm reg}(\langle f_{1},\ldots,f_{m}\rangle_{R})$, it follows that $Z\cap\Lambda^{\rm gcsr}$ contains a non-empty Zariski open set in $Z$ equipped with the induced topology, which implies that [29, Conjecture 4.3.4] is true for the parameter $(n,m,d_{1},\ldots,d_{m})$.
3 Proofs of main results
Let $R=K[x_{1},\ldots,x_{n}]$ be the polynomial ring of $n$ variables over a field $K$ (in any characteristic), and let $I$ be a proper homogeneous ideal of $R$. Let $F=\{f_{1},\ldots,f_{m}\}$ be a subset of $R$ generating $I$, and put $d_{j}:=\deg(f_{j})$ (total degree) for $1\leq j\leq m$.
For a linear transformation $\sigma$ of variables $x_{1},\ldots,x_{n}$ over $K$ represented by a $n\times n$ matrix $P$ over $K$, we denote by $h^{\sigma}:=h(\bm{x}\cdot P)$ the image of $h\in R$ by $\sigma$, where $\bm{x}:=(x_{1},\ldots,x_{n})$.
3.1 Proofs of Theorem 1 (1) and (2)
Let $\prec$ be the graded reverse lexicographic ordering on $R$ with $x_{n}\prec\cdots\prec x_{1}$. We denote by $N$ the number of its projective zeros over $\overline{K}$ of the homogeneous ideal $I\subset R$. Throughout this subsection, assume that $\mathrm{Krull.dim}(R/I)\leq 1$, i.e., $N<\infty$.
Lemma 3.1.1.
If the cardinality of $K$ is greater than $N$, there exist a non-negative integer $d$ and a linear form $\ell\in R$ with $\ell\notin I$ such that the $K$-linear map $(R/I)_{d-1}\to(R/I)_{d}$ defined by the multiplication by $\ell$ is surjective.
Proof.
The assertion is proved in the proof of (i) $\Rightarrow$ (iv) in [11, Theorem 3.3.4] when $K$ is an infinite field: In fact, it suffices for the proof that one can take a linear form $\ell\in R$ that does not vanish at any projective zero over $\overline{K}$ of $I$. Even in the case where $K$ is an finite field $\mathbb{F}_{q}$, if $q$ is greater than the number $N$ of projective zeros over $\overline{K}$ of $I$, we prove that such a linear form always exists.
For each projective zero $\bm{a}=(a_{1}:\cdots:a_{n})\in\mathbb{P}^{n-1}(\overline{K})$ of $I$, we denote by $S_{\bm{a}}$ the set of all linear forms that vanish at $\bm{a}$, say
$S_{\bm{a}}=\{y_{1}x_{1}+\cdots+y_{n}x_{n}\mid y_{i}\in K\mbox{ and }y_{1}a_{1}+\cdots y_{n}a_{n}=0\},$
which is isomorphic, as $K$-linear spaces, to the following $K$-linear space of $K^{n}$:
$S_{\bm{a}}^{\prime}=\{(y_{1},\ldots,y_{n})\in K^{n}\mid a_{1}y_{1}+\cdots a_{n}y_{n}=0\}.$
The dimension of $S_{\bm{a}}^{\prime}$ as a $K$-linear space is at most $n-1$. Indeed, for an arbitrary non-zero entry $a_{i}$ of $\bm{a}$, the vector with $1$ in the $i$-th entry and 00’s elsewhere does not belong to $S_{\bm{a}}^{\prime}$, so that $S_{\bm{a}}^{\prime}$ is a proper subspace of $K^{n}$.
Here, we have
#$\#\left(\bigcup_{\bm{a}\in V_{\overline{K}}(I)}S_{\bm{a}}\right)\leq\sum_{{\bm{a}\in V_{\overline{K}}(I)}}\#S_{\bm{a}}\leq Nq^{n-1}<q^{n},$
whence a desired linear form over $K$ exists. ∎
Proof of Theorem 1 (1):
Note that, by a choice of $\ell$, the Krull dimension of $R/\langle I,\ell\rangle$ is zero. For $d^{\prime}\geq 1$, the $-\ell$ map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ is surjective if and only if $(R/I)_{d^{\prime}}=0$, which follows from the following exact sequence:
${(R/I)_{d^{\prime}-1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\times\ell}$${(R/I)_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(R/\langle I,\ell\rangle)_{d^{\prime}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0.}$
When $d^{\prime}=d$, we have $(R/\langle I,\ell\rangle)_{d}=0$, by the surjectivity of the multiplication map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ by $\ell$. Thus, it follows that $(R/\langle I,\ell\rangle)_{d^{\prime}}=0$ for any $d^{\prime}\geq d$, from which the $-\ell$ map $(R/I)_{d^{\prime}-1}\to(R/I)_{d^{\prime}}$ is surjective for any $d^{\prime}$ with $d^{\prime}\geq d$. Note that the minimal $d$ for which the map is surjective is equal to $d_{\rm reg}(\langle I,\ell\rangle_{R})$.
We first note that $\widetilde{d}_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I^{\tau})$ and ${d}_{\rm reg}(\langle I,\ell\rangle_{R^{\prime}})={d}_{\rm reg}(\langle I^{\tau},\ell^{\tau}\rangle_{R^{\prime}})$ for any invertible linear transformation $\tau$ of variables. There exists a $k\in\{1,\ldots,n\}$ such that $x_{k}$-coefficient of $\ell$ is not zero. We may also assume that the $x_{k}$-coefficient is $1$, without loss of generality. Let $\sigma_{1}$ be the linear transformation of variables $x_{1},\ldots,x_{n}$ defined by a permutation sending $x_{k}$ to $x_{n}$, and write $\ell^{\sigma_{1}}=a_{0}x_{0}+a_{1}x_{1}+\cdots+a_{n}x_{n}$ for $a_{i}\in K$ with $a_{n}=1$. Then, let $\sigma_{2}$ be a linear transformation of variables $x_{1},\ldots,x_{n}$ represented by the matrix
$P:=\begin{pmatrix}1&&&-a_{1}\\ &\ddots&&\vdots\\ &&1&-a_{n-1}\\ &&&1\\ \end{pmatrix},$
where blank entries are 00’s. Namely, for $h\in R$, we define
$h^{\sigma_{2}}:=h(\bm{x}\cdot P)=h(x_{1},\ldots,x_{n-1},x_{n}-a_{1}x_{1}-\cdots-a_{n-1}x_{n-1})$
with $\bm{x}:=(x_{1},\ldots,x_{n})$. Then, the invertible transformation $\sigma:=\sigma_{2}\circ\sigma_{1}$ of variables $x_{1},\ldots,x_{n}$ sends $\ell$ to $x_{n}$, and we have the equalities $D^{\prime}:=\widetilde{d}_{\rm reg}(I)=\widetilde{d}_{\rm reg}(I^{\sigma})$ and $D:=d_{\rm reg}(\langle I,\ell\rangle_{R})=d_{\rm reg}(\langle I^{\sigma},x_{n}\rangle_{R})$. The assertion follows from Theorem 2.1.9.∎
In the following, we shall prove Theorem 1 (2), for which we first prove the following easy lemma in the case where $m=n-1$:
Lemma 3.1.2.
With notation as above, if $m=n-1$ and if the sequence $(f_{1},\ldots,f_{m},\ell)$ is cryptographic semi-regular (equivalently, it is regular since $m+1=n$), then we have $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(\langle I,\ell\rangle_{R})-1$, and hence
$\mathrm{max.GB.deg}_{\prec}(I^{\sigma})\leq\mathrm{max}\{d_{\rm reg}(\langle I,\ell\rangle_{R}),\widetilde{d}_{\rm reg}(I)\}=d_{\rm reg}(\langle I,\ell\rangle_{R})=\sum_{j=1}^{n-1}(d_{j}-1)+1.$
Moreover, $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular (equivalently, it is regular since $m<n$).
Proof.
Recall that the the regularity of a sequence of homogeneous polynomials is equivalent to the semi-regularity if the number of polynomials is less than or equal to the number of variables. For the sequence $(f_{1},\ldots,f_{m},\ell)$, the number of polynomials and that of variables are both $n$, whence it is regular. Therefore, $(f_{1},\ldots,f_{m})$ is also regular, and
${\rm HS}_{R/I}(z)={\rm HS}_{R/\langle I,\ell\rangle}(z)\cdot\frac{1}{1-z}={\rm HS}_{R/\langle I,\ell\rangle}(z)\cdot(1+z+z^{2}+\cdots)$
with
${\rm HS}_{R/\langle I,\ell\rangle}(z)=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})\cdot(1-z)}{(1-z)^{n}}=\frac{\prod_{j=1}^{n-1}(1-z^{d_{j}})}{(1-z)^{n-1}}=\prod_{j=1}^{n-1}(1+z+\cdots+z^{d_{j}-1}),$
which is a polynomial. Here, we have $\mathrm{HF}_{R/I}(d)=\sum_{i=0}^{d}\mathrm{HF}_{R/\langle I,\ell\rangle}(i)$ for any $d$. Since $\mathrm{HF}_{R/\langle I,\ell\rangle}(d)=0$ (resp. $\mathrm{HF}_{R/\langle I,\ell\rangle}(d)\neq 0$) for $d\geq d_{\rm reg}(\langle I,\ell\rangle_{R})$ (resp. $d\leq d_{\rm reg}(\langle I,\ell\rangle_{R})-1$), we have $\widetilde{d}_{\rm reg}(I)=d_{\rm reg}(\langle I,\ell\rangle_{R})-1$, as desired. ∎
Proof of Theorem 1 (2):
Suppose that $(f_{1},\ldots,f_{m})$ is generalized cryptographic semi-regular. Then, applying Lemma 2.2.8, we have the inequality (2.2.3), the right hand side of which is equal to $D_{\rm new}$ for $m\geq n$. It also follows from Lemma 3.1.2 that $\mathrm{max.GB.deg}_{\prec}(I)$ is upper-bounded by $\sum_{j=1}^{n-1}(d_{j}-1)+1$, which is equal to $D_{\rm new}$ for $m=n-1$, as desired.∎
3.2 Proof of Proposition 1
As in the previous subsection, let $\prec$ be the graded reverse lexicographic ordering on $R=K[x_{1},\ldots,x_{n}]$ with $x_{n}\prec\cdots\prec x_{1}$. We also denote the restriction of $\prec$ on $K[x_{1},\ldots,x_{n-1}]$ by the same notation $\prec$. In the following, we shall prove Proposition 1, namely the inequality (1.0.1) becomes an equality if $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic. Here, a weakly reverse lexicographic ideal is a monomial ideal $J$ such that if $x^{\alpha}$ is one of the minimal generators of $J$ then every monomial of the same degree which preceeds $x^{\alpha}$ must belong to $J$ as well (see also e.g., [34, Section 4] for the definition).
As in [23] and [24] (see also [28] and [29]), we consider the elimination ideal
$I^{\prime}:=\langle I,x_{n}\rangle_{R}\cap K[x_{1},\ldots,x_{n-1}]=\langle(f_{1})|_{x_{n}=0},\ldots,(f_{m})|_{x_{n}=0}\rangle_{K[x_{1},\ldots,x_{n-1}]}.$
Since $R/\langle I,x_{n}\rangle$ and $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ are isomorphic as graded rings, they have the same Hilbert series, and $d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$. Put $D:=d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$. By comparing Hilbert series, it can be also easily checked that, the sequence $(f_{1},\ldots,f_{m},x_{n})\in R^{m+1}$ is cryptographic semi-regular if and only if $(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})\in K[x_{1},\ldots,x_{n-1}]^{m}$ is cryptographic semi-regular, equivalently these sequences are both $D$-regular.
Let $G$ (resp. $G^{\prime}$) be the reduced Gröbner basis of $I$ (resp. $I^{\prime}$) with respect to $\prec$. Note that, since $\langle I,x_{n}\rangle_{R}=\langle I^{\prime},x_{n}\rangle_{R}$, the reduced Gröbner basis of $\langle I,x_{n}\rangle_{R}$ is given by $G^{\prime}\cup\{x_{n}\}$, whence $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=\mathrm{max.GB.deg}_{\prec}(\langle I,x_{n}\rangle_{R})$.
To prove Proposition 1, we start with extending some results in [28] to our case.
Lemma 3.2.1 (cf. [28, Corollary 2]).
With notation as above, we assume that $(f_{1},\ldots,f_{m},x_{n})\in R^{m+1}$ is cryptographic semi-regular (i.e., $(f_{1}^{\rm top},\ldots,f_{m}^{\rm top})\in K[x_{1},\ldots,x_{n-1}]^{m}$ is cryptographic semi-regular), namely both the sequences are $D$-regular with $D:=d_{\rm reg}(I^{\prime})=d_{\rm reg}(\langle I,x_{n}\rangle_{R})$. Then, we have
$(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}=(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d}$ (3.2.1)
for any $d$ with $d<D$.
Proof.
We first prove $(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}\supset(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d}$ for any non-negative integer $d$. For this, it suffices to show that $\mathrm{LM}_{\prec}(f)\in\mathrm{LM}_{\prec}(I)$ for any $f\in I_{d}^{\prime}\smallsetminus\{0\}$. We may assume that $f$ is homogeneous, and then it is written as $f=g+x_{n}h$ for some $g\in I_{d}$ and $h\in R_{d-1}$. Here, $g$ is not divisible by $x_{n}$, since otherwise $f\notin K[x_{1},\ldots,x_{n-1}]$, a contradiction. Therefore, we have $\mathrm{LM}_{\prec}(f)=\mathrm{LM}_{\prec}(g)$ with $g\in I$, as desired.
Since $(f_{1},\ldots,f_{m},x_{n})$ is $D$-regular, it follows from [14, Proposition 2 a)] (cf. [28, Corollary 1]), its sub-sequence $(f_{1},\ldots,f_{m})$ is also $D$-regular. Hence, we have
${\rm HS}_{R/I}(z)$ $\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})}{(1-z)^{n}}\equiv\frac{\prod_{j=1}^{m}(1-z^{d_{j}})\cdot(1-z)}{(1-z)^{n}}\cdot\frac{1}{1-z}$
$\equiv{\rm HS}_{R/\langle I,x_{n}\rangle}(z)\cdot\frac{1}{1-z}\equiv{\rm HS}_{R/\langle I,x_{n}\rangle}(z)\cdot(1+z+z^{2}+\cdots)$
$\equiv{\rm HS}_{K[x_{1},\ldots,x_{n-1}]/I^{\prime}}(z)\cdot(1+z+z^{2}+\cdots)\pmod{z^{D}}.$
Thus, for any $d$ with $d<D=d_{\rm reg}(I^{\prime})$, one has ${\rm HF}_{R/I}(d)=\sum_{i=0}^{d}\mathrm{HF}_{K[x_{1},\ldots,x_{n-1}]/I^{\prime}}(i)$, so that
$\dim_{K}R_{d}-\dim_{K}I_{d}=\sum_{i=0}^{d}(\dim_{K}K[x_{1},\ldots,x_{n-1}]_{i}-\dim_{K}I^{\prime}_{i}).$ (3.2.2)
Since $(JR)_{d}=x_{n}^{d}J_{0}\oplus x_{n}^{d-1}J_{1}\oplus\cdots\oplus x_{n}J_{d-1}\oplus J_{d}$ for any of $J=I^{\prime}$ and $J=K[x_{1},\ldots,x_{n-1}]$ (note that this holds also for $d\geq D$), the right hand side of (3.2.2) is equal to $\dim_{K}R_{d}-\dim_{K}(I^{\prime}R)_{d}$, so that $\dim_{K}(I^{\prime}R)_{d}=\dim_{K}I_{d}$. Here, it follows from [22, Theorem 5.2.6] that $\dim_{K}(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d}=\dim_{K}(\langle\mathrm{LM}_{\prec}(I^{\prime}R)\rangle_{R})_{d}$. Since $\langle\mathrm{LM}_{\prec}(I^{\prime}R)\rangle_{R}=\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R}$, we finally obtain (3.2.1). ∎
Lemma 3.2.2 (cf. [28, Lemma 2]).
Under the same setting as in Lemma 3.2.1, for each degree $d<D$, we have
${\rm LM}_{\prec}(G)_{d}={\rm LM}_{\prec}(G^{\prime})_{d}.$ (3.2.3)
Proof.
The equality (3.2.3) can be proved by the induction on $d$, as follows. The case $d=0$ is trivial. Suppose that (3.2.3) holds for any $d$ with $10\leq d<D-$ 1.
Let $t$ be any monomial in ${\rm LM}_{\prec}(G)_{d+1}$, so that $t=\mathrm{LM}_{\prec}(g)$ for some $g\in G$. By $d+1<D$, it follows from Lemma 3.2.1 that
$(\langle\mathrm{LM}_{\prec}(I)\rangle_{R})_{d+1}=(\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{R})_{d+1},$
whence ${\rm LM}_{\prec}(g)$ is divided by ${\rm LM}_{\prec}(g^{\prime})$ for some $g^{\prime}\in G^{\prime}$. Since $G$ is reduced, the induction hypothesis implies that $\mathrm{deg}(g^{\prime})=d+1$, and thus $\mathrm{LM}_{\prec}(g)=\mathrm{LM}_{\prec}(g^{\prime})$. Therefore, we obtain ${\rm LM}_{\prec}(G)_{d+1}\subset{\rm LM}_{\prec}(G^{\prime})_{d+1}$. The same argument shows that ${\rm LM}_{\prec}(G)_{d+1}\supset{\rm LM}_{\prec}(G^{\prime})_{d+1}$, as desired. ∎
Lemma 3.2.3.
Under the same setting as in Lemma 3.2.1, we also suppose that $R/\langle I,x_{n}\rangle$ is Artinian (equivalently $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ is Artinian), namely $D<\infty$. Furthermore, we assume that $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic. Then we have $\mathrm{max.GB.deg}_{\prec}(I)\geq D$ and $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=D$.
Proof.
Put $d:=\mathrm{max.GB.deg}_{\prec}(I)$. Assume for a contradiction that $d<D$. Since $\mathrm{Krull.dim}(R/\langle I,x_{n}\rangle)=0$, it follows from Lemma 2.1.5 that there is an element $g$ in $G$ with ${\rm LM}_{\prec}(g)=x_{n-1}^{d^{\prime}}$ for some $d^{\prime}\leq d$. Then, any monomial $t\in R_{d^{\prime}}$ with $t\succeq x_{n-1}^{d^{\prime}}$ (namely $t$ is any monomial in $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}$) belongs to $\langle{\rm LM}_{\prec}(I)\rangle_{R}=\langle{\rm LM}_{\prec}(G)\rangle_{R}$ by the weakly reverse lexicographicness. Here, by $d<D$ together with (3.2.3), we have $\langle{\rm LM}_{\prec}(G)\rangle_{R}=\langle{\rm LM}_{\prec}(G^{\prime})_{\leq d}\rangle_{R}$. This implies that $t$ also belongs to $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$, and hence $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}/I^{\prime}_{d^{\prime}}=0$. Thus, we have $d\geq d^{\prime}\geq d_{\rm reg}(I^{\prime})$, a contradiction.
By the same argument as above, we can show that $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=d_{\rm reg}(I^{\prime})$ as follows. Note that $\mathrm{max.GB.deg}_{\prec}(I^{\prime})\leq d_{\rm reg}(I^{\prime})$ follows from a general fact written in the second paragraph of Remark 3.2.4 below. Thus, we assume to the contrary that $d:=\mathrm{max.GB.deg}_{\prec}(I^{\prime})<D:=d_{\rm reg}(I^{\prime})$. Then, it follows from (3.2.3) that ${\rm LM}_{\prec}(G^{\prime})={\rm LM}_{\prec}(G)_{\leq d}$ and the ideal $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ has the weak reverse lexicographicness up to $d$. Since there is an element $g$ in $G^{\prime}$ with ${\rm LM}_{\prec}(g)=x_{n-1}^{d^{\prime}}$ for some $d^{\prime}\leq d$, any monomial in $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}$ belongs to $\langle{\rm LM}_{\prec}(G^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ and so $K[x_{1},\ldots,x_{n-1}]_{d^{\prime}}/I^{\prime}_{d^{\prime}}=0$, which implies $d\geq d^{\prime}\geq D$, a contradiction. ∎
Remark 3.2.4.
In Lemma 3.2.3, when $R/\langle I,x_{n}\rangle$ is Artinian (or equivalently $K[x_{1},\ldots,x_{n-1}]/I^{\prime}$ is Artinian), we can easily prove $\mathrm{max.GB.deg}_{\prec}(I^{\prime})=D$ if $\langle\mathrm{LM}_{\prec}(I^{\prime})\rangle_{K[x_{1},\ldots,x_{n-1}]}$ is a weakly reverse lexicographic ideal, assuming neither that $(f_{1},\ldots,f_{m},x_{n})$ is cryptographic semi-regular nor that $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}$ is weakly reverse lexicographic.
In general, for any homogeneous ideal $J\subset R$ with $\mathrm{Krull.dim}(R/J)=0$, we have $\mathrm{max.GB.deg}_{\prec}(J)\leq\mathrm{reg}(J)=d_{\rm reg}(J)$ by [4, Corollary 2.5] (see also [23, Proposition 4.15]), and it is straightforward that the equality holds if $\langle\mathrm{LM}_{\prec}(J)\rangle_{R}$ is a weakly reverse lexicographic.
Proof of Proposition 1:
Put $D:=d_{\rm reg}(\langle I,x_{n}\rangle_{R})=d_{\rm reg}(I^{\prime})$ and $D^{\prime}:=\widetilde{d}_{\rm reg}(I)$. From our assumption that $R/\langle I,x_{n}\rangle$ is Artinian, we have the inequality (1.0.1), i.e., $\mathrm{max.GB.deg}(I)\leq\mathrm{max}\{D,D^{\prime}\}$. The case where $\mathrm{Krull.dim}(R/I)=0$ is clear from Remark 3.2.4, so we assume that $\mathrm{Krull.dim}(R/I)=1$. Note that, for the reduced Gröbner basis $G$ of $I$, the elements of $\mathrm{LM}_{\prec}(G)$ are the minimal generators of the monomial ideal $\langle\mathrm{LM}_{\prec}(I)\rangle_{R}=\langle\mathrm{LM}_{\prec}(G)\rangle_{R}$.
First we consider the case where $D^{\prime}\geq D$, so that $\max\{D,D^{\prime}\}=D^{\prime}$. Put $d:=\mathrm{max.GB.deg}_{\prec}(I)$. Then, it follows from Macaulay’s basis theorem (cf. [26, Theorem 1.5.7]) that $R_{d}/I_{d}$ has a $K$-linear basis of the form
$\{t\in R_{d}:\mbox{$t$ is a monomial and }{t\notin\langle{\rm LM}_{\prec}(I)\rangle_{R}=\langle\mathrm{LM}_{\prec}(G)\rangle_{R}}\},$
which is called the standard monomial basis. We set $r:=\dim_{K}R_{d}/I_{d}$, and write this standard monomial basis as $\{t_{1},\ldots,t_{r}\}$. In the following, we prove that $\{t_{1}x_{n}^{s},\ldots,t_{r}x_{n}^{s}\}$ is a $K$-linear basis of $R_{d+s}/I_{d+s}$ for any $s$ with $s\geq 1$, from which we have $\dim_{K}R_{d}/I_{d}=\dim_{K}R_{d+s}/I_{d+s}$ for any positive integer $s$, so that $d\geq D^{\prime}$ and therefore $d=D^{\prime}$.
By Lemma 3.2.3, we have $d\geq D$, and thus it can be shown similarly to Proof of Theorem 1 (1) that the multiplication-by-$x_{n}^{s}$ map from $R_{d}/I_{d}$ to $R_{d+s}/I_{d+s}$ is surjective (since $R/\langle I,x_{n}\rangle$ is Artinian). Therefore $B_{s}:=\{t_{1}x_{n}^{s},\ldots,t_{r}x_{n}^{s}\}$ generates $R_{d+s}/I_{d+s}$. Suppose to contrary that $B_{s}$ is not a basis. In this case, there exists an $i$ such that $t_{i}x_{n}^{s}$ is divisible by $\mathrm{LM}_{\prec}(g)$ for some $g\in G$. Indeed, since the elements of $B_{s}$ are linearly dependent over $K$, we obtain $f:=a_{1}t_{1}x_{n}^{s}+\cdots+a_{r}t_{r}x_{n}^{s}\in I_{d+s}$ for some $(a_{1},\ldots,a_{r})\in K^{r}\smallsetminus\{(0,\ldots,0)\}$. Thus, for the least $i$ with $a_{i}\neq 0$, the leading monomial of $f$ is $t_{i}x_{n}^{s}$, which is divisible by $\mathrm{LM}_{\prec}(g)$ for some $g\in G$. Putting $u_{i}=\mathrm{GCD}(t_{i},\mathrm{LM}_{\prec}(g))$ and $s_{i}=t_{i}/u_{i}$, we have $u_{i}s_{i}x_{n}^{s}=t_{i}x_{n}^{s}$. Since $t_{i}=s_{i}u_{i}$ is not divisible by $\mathrm{LM}_{\prec}(g)$, we can write $u_{i}x_{n}^{s^{\prime}}=\mathrm{LM}_{\prec}(g)$ for some $s^{\prime}\leq s$. (We note that $\deg s_{i}=d-\deg u_{i}\geq\deg{\rm LM}_{\prec}(g)-\deg u_{i}=s^{\prime}$.) Note that $s^{\prime}\geq 1$ since otherwise $t_{i}$ is divisible by $\mathrm{LM}_{\prec}(g)$.
Take an arbitrary monomial $s^{\prime}_{i}$ such that $\deg s^{\prime}_{i}=s^{\prime}$ and $s_{i}^{\prime}$ divides $s_{i}$. Then, by the weakly reverse lexicographicness, as ${\rm LM}_{\prec}(g)=u_{i}x_{n}^{s^{\prime}}\preceq u_{i}s^{\prime}_{i}$, the monomial $s^{\prime}_{i}u_{i}$ should belong to $\langle{\rm LM}_{\prec}(G)\rangle_{R}$. Moreover, since $s^{\prime}_{i}u_{i}$ divides $t_{i}=s_{i}u_{i}$, the monomial $t_{i}=s_{i}u_{i}$ also belongs to $\langle{\rm LM}_{\prec}(G)\rangle_{R}$, which is a contradiction.
Finally, consider the remaining case, i.e., $D>D^{\prime}$, so that $\max\{D,D^{\prime}\}=D$. In this case, it follows from Lemma 3.2.3 that $d\geq D$, whence $d=D$.∎
3.3 Proof of Theorem 1 (3): Complexity estimation
We shall prove (3) of Theorem 1. For this, we estimate the complexity of computing a Gröbner basis of a homogeneous ideal with Macaulay matrices.
Let $F=\{f_{1},\ldots,f_{m}\}\subset R$ be a set of homogeneous polynomials in $R\smallsetminus K$, put $d_{j}:=\deg(f_{j})$ for $1\leq j\leq m$, and let $\prec$ be an arbitrary graded monomial ordering on $R$. The complexity of the Gröbner basis computation for $F$ is estimated as that of computing the reduced row echelon form (RREF) of the degree-$D$ Macaulay matrix, denoted by $M_{\leq D}(F)$, where $D$ is an arbitrary upper-bound on $\mathrm{max.GB.deg}_{\prec}(F)$. Here, the degree-$D$ Macaulay matrix $M_{\leq D}(F)$ (in homogeneous case) is a block matrix of the form
$M_{\leq D}(F)=\begin{pmatrix}M_{D}(F)&&\\ &\ddots&\\ &&M_{d_{0}}(F)\end{pmatrix}$
with $d_{0}:=\min\{\deg(f_{j}):1\leq j\leq m\}$, where each $M_{d}(F)$ is a homogeneous Macaulay matrix defined as follows: For each $d$ with $d\geq d_{0}$, the degree-$d$ homogeneous Macaulay matrix $M_{d}(F)$ of $F$ has columns indexed by the terms of $R_{d}$ sorted, from left to right, according to the chosen order $\prec$. The rows of $M_{d}(F)$ are indexed by the polynomials $m_{i,j}f_{j}$, where $m_{i,j}\in R$ is a term such that $\deg(m_{i,j}f_{j})=d$.
Theorem 3.3.1.
With notation as above, we have the following:
  1. (1)
    [2, Proposition 1] The RREF of $M_{\leq D}(F)$ is computed in
    $O\left(mD\binom{n+D-1}{D}^{\omega}\right).$ (3.3.1)
    Hence, if $\prec$ is a graded order, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.3.2) for $D=\mathrm{max.GB.deg}_{\prec}(F)$.
  2. (2)
    Let $D$ be a non-negative integer such that $D=O(n)$. Then, the RREF of $M_{\leq D}(F)$ is computed in
    $O\left(m\binom{n+D-1}{D}^{\omega}\right).$ (3.3.2)
    Hence, if $\prec$ is a graded order, and if $\mathrm{max.GB.deg}_{\prec}(F)=D=O(n)$, then the complexity of the Gröbner basis computation for $F$ is upper-bounded by (3.3.2).
Proof.
It suffices to estimate the complexity of computing the RREFs of the homogeneous Macaulay matrices $M_{D}(F),\ldots,M_{d_{0}}(F)$. For each $d$ with $d_{0}\leq d\leq D$, the number of rows (resp. columns) in $M_{d}(F)$ is $\sum_{1\leq i\leq m,\ d\geq d_{i}}\binom{n-1+d-d_{i}}{d-d_{i}}$ (resp. $\binom{n-1+d}{d}$). It follows from $d-d_{i}\leq d-1$ that
$\sum_{1\leq i\leq m,\ d\geq d_{i}}\binom{n-1+d-d_{i}}{d-d_{i}}\leq\sum_{i=1}^{m}\binom{n-1+d-1}{d-1}=m\binom{n-1+d-1}{d-1},$
whence the number of rows in $M_{d}(F)$ is upper-bounded by $m\binom{n-1+d-1}{d-1}$. Therefore, the complexity of the row reductions on $M_{D}(F),\ldots,M_{d_{0}}(F)$ is upper-bounded by
$\sum_{d=d_{0}}^{D}m\binom{n-1+d-1}{d-1}\binom{n-1+d}{d}^{\omega-1}$ $\leq m\sum_{d=1}^{D}\binom{n-1+d-1}{d-1}\binom{n-1+d}{d}^{\omega-1}$
$\leq m\sum_{d=1}^{D}\binom{n-1+d}{d}^{\omega},$ (3.3.3)
where we used a fact that the reduced row echelon form of a $k\times\ell$ matrix $A$ over a field $K$ can be computed in $O(k\ell^{\omega-1})$, see Remark 3.3.2 below for its proof. Clearly (3.3.3) is upper-bounded by $mD\binom{n+D-1}{D}^{\omega}$, we obtain the estimation (3.3.1) in (1).
In the following, we shall improve the estimation (3.3.1) in (1) to obtain (3.3.2) in (2), assuming that $D=O(n)$. Here, for each $d$ with $11\leq d\leq D-$ 1, we have
$\binom{n-1+d}{d}=\frac{(n-1+d)!}{(n-1)!d!}=\binom{n-1+d+1}{d+1}\times\frac{d+1}{n+d}.$
Repeating this procedure, we obtain
$\binom{n-1+d+1}{d+1}\times\frac{d+1}{n+d}=\binom{n-1+d+2}{d+2}\times\frac{d+2}{n+d+1}\times\frac{d+1}{n+d}$
$=\cdots=\binom{n-1+D}{D}\times\frac{D}{n+D-1}\times\cdots\times\frac{d+2}{n+d+1}\times\frac{d+1}{n+d}.$
It is clear that $\frac{d^{\prime}+1}{n+d^{\prime}}=1-\frac{n-1}{n+d^{\prime}}$ is monotonically increasing with respect to $d^{\prime}$. Putting $c:=(D-1)/n$, we have $D=cn+1$, so that
$\frac{d^{\prime}+1}{n+d^{\prime}}\leq\frac{D}{n+D-1}=\frac{cn+1}{(c+1)n}=\frac{c+\frac{1}{n}}{c+1}$
for any $d^{\prime}$ with $d^{\prime}\leq D-1$. As we consider $n\to\infty$, we may suppose that
$\frac{d^{\prime}+1}{n+d^{\prime}}\leq\frac{c+\frac{1}{n}}{c+1}<\frac{c+\varepsilon}{c+1}<1$
for some constant $\varepsilon$ with $10<\varepsilon<$ 1. Therefore, we obtain
$\binom{n-1+D}{D}\times\frac{D}{n+D-1}\times\cdots\times\frac{d+2}{n+d+1}\times\frac{d+1}{n+d}<\binom{n-1+D}{D}\left(\frac{c+\varepsilon}{c+1}\right)^{D-d},$
whence
$\sum_{d=1}^{D}\binom{n-1+d}{d}^{\omega}$ $<\binom{n-1+D}{D}^{\omega}\sum_{k=0}^{D-1}\left(\frac{c+\varepsilon}{c+1}\right)^{k\omega}$
$=\binom{n-1+D}{D}^{\omega}\times\frac{1-\left(\frac{c+\varepsilon}{c+1}\right)^{D\omega}}{1-\left(\frac{c+\varepsilon}{c+1}\right)^{\omega}}$
$<\binom{n-1+D}{D}^{\omega}\times\frac{1}{1-\left(\frac{c+\varepsilon}{c+1}\right)^{\omega}},$
where we put $k:=D-d$. It follows from $D=O(n)$ that $c$ is upper-bounded by a constant. By $10<\varepsilon<$ 1, it is clear that $\frac{c+\varepsilon}{c+1}$ is also upper-bounded by a positive constant smaller than $1$, whence
$\frac{1}{1-\left(\frac{c+\varepsilon}{c+1}\right)^{\omega}}=O(1)$
as $n\to\infty$. We have proved the assertion (2). ∎
Remark 3.3.2.
In the proof of Theorem 3.3.1, we used a fact that the reduced row echelon form of a $k\times\ell$ matrix $A$ over a field $K$ can be computed in $O(k\ell^{\omega-1})$, where $k$ can be greater than $\ell$. This fact is proved by considering the following procedures to compute the reduced row echelon form of $A$:
  1. (0)
    Write $k=\ell q+r$ for $q,r\in\mathbb{Z}$ with $0\leq r<\ell$.
  2. (1)
    Choose and remove $2\ell$ rows from $A$, and let $A^{\prime}$ be a $2\ell\times\ell$ matrix whose rows are the chosen $2\ell$ rows.
  3. (2)
    Compute the reduced row echelon form of $A^{\prime}$: Inserting $\ell$ zeros to the end of each row of $A^{\prime}$, construct a $2\ell\times 2\ell$ square matrix, and compute its reduced row echelon form.
  4. (3)
    As a result of (2), we obtain more than or equal to $\ell$ zero row vectors. We add the nonzero (linearly independent) row vectors obtained to $A$.
  5. (4)
    Go back to (1).
Repeating the procedures from (1) to (4) at most $q$ times, we obtain the reduced row echelon form $B$ of the initial $A$. Since the complexity of (2) is $(2\ell)^{\omega}$, the total complexity to obtain $B$ is $q\cdot(2\ell)^{\omega}\leq(k/\ell)\cdot 2^{\omega}\ell^{\omega}=O(k\ell^{\omega-1})$, as desired.
Proof of Theorem 1 (3):
For the graded reverse lexicographic ordering $\prec$ on $R$ with $x_{n}\prec\cdots\prec x_{1}$, it follows from Theorem 1 (2) that $D_{\rm new}$ gives an upper bound on $\mathrm{max.GB.deg}_{\prec}(F)$. Therefore, a Gröbner basis of $\langle F\rangle_{R}$ is obtained by computing the RREF of $M_{\leq D}(F)$ for $D=D^{(n,m)}$. Here, the value $D^{(n,m)}$ of (1.0.4) is maximized at $m=n$: In this case, $D^{(n,m)}$ is equal to the Lazard’s bound $\sum_{j=1}^{n}(d_{j}-1)+1$, which is $O(n)$ for fixed $d_{1},\ldots,d_{m}$. Therefore, it follows from Theorem 3.3.1 (2) that the RREF of $M_{\leq D}(F)$ for $D=D^{(n,m)}$ is computed in the complexity (1.0.5), as desired.∎
Acknowledgement
This work was supported by JSPS Grant-in-Aid for Young Scientists 23K12949, JSPS Grant-in-Aid for Scientific Research (C) 21K03377, and JST CREST Grant Number JPMJCR2113.